OpenStudy (anonymous):

Find the solution to the equations. 3x - y = -4 x + y = 0 (0, 0) (-1, 1) (1, -1)

HanAkoSolo (jamierox4ev3r):

you can solve this problem using substitution. first, solve the second equation for y. basically, we want to isolate the "y" variable here

HanAkoSolo (jamierox4ev3r):

@Sydney77 let me know if you understand :)

OpenStudy (anonymous):

Simple

OpenStudy (ageta):

I will try to solve your system of equations. 3x−y=−4;x+y=0 Step: Solve 3x−y=−4 for y: 3x−y+−3x=−4+−3x(Add -3x to both sides) −y=−3x−4 (Divide both sides by -1) y=3x+4 Step: Substitute (3x+4) for y in x+y=0: x+y=0 x+3x+4=0 4x+4=0(Simplify both sides of the equation) 4x+4+−4=0+−4(Add -4 to both sides) (Divide both sides by 4) x=−1 Step: Substitute (−1) for x in y=3x+4: y=3x+4 y=(3)(−1)+4 y=1(Simplify both sides of the equation) Answer: y=1 and x=−1